Exercise 13-19 (Part Level Submission)

Presented below is information related to Riverbed Inc.

RIVERBED INC.
BALANCE SHEET
DECEMBER 31, 2017

Cash $44,700 Notes payable (short-term) $50,400
Receivables $109,000 Accounts payable 31,900
   Less: Allowance 15,000 94,000 Accrued liabilities 5,000
Inventory 171,700 Common stock (par $5) 261,100
Prepaid insurance 8,100 Retained earnings 141,500
Land 20,000
Equipment (net) 151,400   
$489,900 $489,900

RIVERBED INC.
INCOME STATEMENT
FOR THE YEAR ENDED DECEMBER 31, 2017

Sales revenue $1,401,100
Cost of goods sold
   Inventory, Jan. 1, 2017 $200,200
   Purchases 792,700
   Cost of goods available for sale 992,900
   Inventory, Dec. 31, 2017 (171,700 )
   Cost of goods sold 821,200
Gross profit on sales 579,900
Operating expenses 170,200
Net income $409,700

Collapse question part

(a)

Compute the following ratios or relationships of Riverbed Inc. Assume that the ending account balances are representative unless the information provided indicates differently. (Round answers to 2 decimal places, e.g. 1.59 or 45.87%.)
(1) Current ratio

times
(2) Inventory turnover

times
(3) Accounts receivable turnover

times
(4) Earnings per share $

(5) Profit margin on sales

%
(6) Return on assets on December 31, 2017

%
Answer #1

1) Current ratio = Current assets/Current liabilities = 318500/87300 = 3.65:1

2) Inventory turnover = 821200/185950 = 4.42 Times

3) Account receivable turnover = 1401100/94000 = 14.91 Times

4) Earning per share = 409700/52220 = 7.85 per share

5) Profit margin = 409700/1401100 = 29.24%

6) Return on assets = 409700/489900 = 83.63%